OpenStudy (anonymous):

-2x-4=-4x and show me the steps thxs

OpenStudy (smurfy14):

x=2

OpenStudy (anonymous):

-2x-4=-4x -4=-4x+2x -4=-2x -4/2=x -1=x

OpenStudy (smurfy14):

-4=-2x (subtracted) divide by -2 so.. x=2

OpenStudy (anonymous):

sorry its x=-2